OpenStudy (anonymous):

put in exponetial form √3:(√(2))

OpenStudy (anonymous):

\[\sqrt[3]{\sqrt{2}}\]

OpenStudy (anonymous):

(2^(1/2)^(1/3) = 2^(1/6)